# Towards Fully FP8 GEMM LLM Training at Scale

### v1

@@ 1 Introduction
Thus, the development of efficient training techniques has become increasingly essential. One of the main research avenues for efficiency is the use of lower-precision number formats to accelerate training on appropriate hardware accelerators. Recently, the use of 8-bit floating-point (FP8) formats has shown promising results [3, 6, 22]. However, the widespread adoption of current approaches is still limited due to suboptimal throughput benefits. One cause of slowdowns is the use of higher precision in those General Matrix Multiplications (GEMMs) which are most sensitive, such as attention score computation, while another issue is the overhead caused by more granular FP8 scaling strategies. One of the key challenges in 8-bit LLM training originates from the relatively narrow dynamic range offered by FP8 formats and thus higher risk of underflows and overflows, especially with the prevalence of large outlier features observed in the LLM’s neural activations during training [4, 28, 9]. To mitigate this issue, modern FP8 training recipes utilise various scaling techniques before casting from higher-precision formats—typically BF16 [13] for activations—to FP8 formats used in matrix multiplications. These scaling approaches help maximize the effective use of FP8’s limited dynamic range, reducing the risk of underflows and overflows.
@@
Recent work has introduced promising FP8 training recipes by employing multiple scaling factors per single tensor [3], allowing for a finer and more precise casting to lower precision. Yet, this comes with an efficiency overhead, diminishing the large gains initially expected from using FP8. Another strategy involves adjusting the standard SwiGLU-based transformer architecture [26] to prevent emergent outliers from occurring [6]. This area of optimization remains underexplored, as most works focus on FP8 GEMMs within the linear projections within the transformer, while maintaining higher precision for other GEMMs, namely those involved in the dot product attention mechanism. We refer to such training strategies simply as FP8 training. We label the approach of also including attention computation as FP8DPA.
@@
In this paper, we introduce FOG: the fast and outlier-guarded set of LLM architectures specifically designed to mitigate large activation outliers and enable efficient large-scale FP8 training with low-overhead scaling strategies. For the first time to our knowledge, this approach enables FP8 GEMMs not only in the linear projection, but also within the attention mechanism of each transformer block, achieving unprecedented throughput improvements of up to 40% in the 8B parameter model scale, while maintaining equivalent downstream performance compared to higher precision baselines. In addition, we present a comprehensive recipe for monitoring, explaining, and predicting training instabilities that might not surface in the early stages of training. This approach provides researchers with greater confidence in the long-term stability of FP8 training recipes, reducing the need for costly, full-scale experiments when using new architectures. Furthermore, we provide an interestingly useful observation about larger models’ tendency to diverge later in training with FP8.
@@
  • We introduce FOG, a novel set of architectures designed to minimise outlier features during training. Our recipe allows stable training with FP8 computation of all GEMMs inside the transformer blocks, surpassing throughput of existing BF16 approaches by 40%. It is a significant step forward towards a purely FP8 matrix multiplication pre-training framework.
  • Our design achieves equivalent quality results to BF16 baselines, while providing a significant speed-up. We empirically attest both performance and stability on various model sizes (0.4B, 1.5B, 8B) and data regimes, up to 14x the Chinchilla optimal data budget [11].
@@ 2 Background
Further examination has shown that not only is SwiGLU an outlier amplifier, but Gated Linear Units (GLUs) in general, as well as pre-normalization layers, suggesting that improper signal propagation is the root cause of outliers [9]. Removing these components and equipping transformers with QK entropy regularization mechanisms such as QK RMS Normalization [10], producing the OP architectures, has been shown to diminish late-stage outliers observed by orders of magnitude, while providing equivalent prediction quality. While OP architectures were shown to be beneficial for post-training quantization, its use for FP8 pre-training remains unexplored. Finally, an alternative to pre-normalization layers are post-normalization layers [16]. Long data regime trainings have confirmed their superiority in terms of training stability with the standard BF16 mixed precision training [19].
@@ 3 FOG: Fast and Outlier-Guarded FP8-suited architectures
Further, the input of the first transformer block is scaled by $\sigma^{-1}$ to maintain unit variance activations at initialization, where $\sigma$ is the chosen standard deviation of the network’s random initialization. Finally, to enhance performance, a learnable normalization block is applied before the residual connections. This takes the form of a LayerScale [29] block, $\mathrm{LayerScale}_{\boldsymbol{\gamma}}(\mathbf{x}):=\boldsymbol{\gamma}\odot\mathbf{x}$, where $\boldsymbol{\gamma}\in\mathbb{R}^{D}$ is a learnable gain vector, or an RMSNorm block, resulting in a post-normalized architecture [16]. In both cases, the learnable gains vector is initialized to $\mathbf{1}/\sqrt{\mathrm{num\_layers}}$ and keeps the residual branch unnormalized, allowing proper signal propagation [9]. Our architecture suite is specified in Table 1.
@@
While the OP architecture already offers several guards to prevent large outliers from occurring, we observed that it remains an impractical choice for FP8DPA training. In Section 5.1 we show that, like all other architectures tested, it suffers a fatal loss divergence early during training. We isolate the two components responsible for OP’s incompatibility with FP8DPA training: the trainable RMSNorm gains vector $\boldsymbol{\gamma}$, and the lack of any normalization. We identify the use of post-normalization as not prone to the outlier tendency pre-normalization networks have.
Figure 2: From OP to FOG-opt step by step. The first run to diverge is OP, while OP with frozen QK RMSNorm is able to survive through the stable phase of training, but it experiences a significant divergence during learning rate cooldown, which starts around 42B tokens in. The converged run adds post-normalization to the previous recipe, resulting in FOG-opt.
Figure 2 ablates the components transitioning from OP to FOG-opt. We can see that freezing the trainable QK RMSNorm gains results in a significantly more stable training. We attribute the early divergence of OP to the fact that uncontrolled QK normalization leads to an explosion of its gains when training in low precision. Note that these gains are generally not weight-decayed. We experimentally observe this explosion, confirm that using $L_{2}$ regularization helps delay the divergence. We finally opted for freezing the gains to a constant value as it is simpler and sufficient, doesn’t compromise performance, and offers a small speedup. Our ablations highlight that a constant value for the gains slightly greater than 1 improves loss. Therefore, to retain its benefit after removing the $\boldsymbol{\gamma}$ gains vector, we increase the standard $s=1/\sqrt{D_{qk}}$ attention softmax temperature–a tiny optimization trick offering equivalent attention score matrix $\mathbf{S}$:
@@
Prior works also favored post-normalization over pre-normalization [19], providing evidence of their better stability in BF16 training. We extend this observation to our FP8 setting and we confirm that learnable LayerScale blocks alone, even with controlled QK regularization, cannot ensure convergence during this last phase. We attribute this late divergence of OP to the fact that LayerScale blocks without normalizations are not enough to handle FP8 outliers, potentially due to the considerable changes in model statistics following the learning rate decay, that are summed up in the residual connections resulting in huge activation outliers for last layers, as highlighted by the increasing pattern of outliers on each transformer block’s output in Figure 3.
@@ 4 Long-term outlier dynamics
Figure 4: Loss and kurtosis training dynamics of 1.5B FOG-max and a pre-normalized GLU-based model trained for a total of 100B tokens. Note the logarithmic kurtosis axis in the transformer block output and FFN activations.
@@
First, these architectures exhibit a sub-linear to logarithmic trend in the long-term growth of QKV outliers, as consistently shown by kurtosis in Figure 4. This behavior supports their robustness to FP8DPA, as it suggests that prohibitively longer training would be required to see a substantial increase in kurtosis. Our extended run is consistent with the hypothesis as it does not exhibit any sign of divergence.
Second, baseline Llama exhibits late divergence during FP8 training (with attention in BF16), which has been attributed solely to the quadratic behavior of its gated activation function—emerging when weights become sufficiently aligned late in training [6]. In our extended 420-billion-token run using the FOG-max architecture, we employ the inherently quadratic xIELU activation function [12] and observe smooth training with kurtosis levels orders of magnitude lower than those of baseline Llama. This challenges the completeness of the prior explanation. From another perspective, maintaining very low kurtosis throughout training enables the stable use of quadratic activations. This is particularly interesting given that such activations are known to produce linear gradients, which benefit the backward pass—likely contributing to FOG-max’s superior performance over GeLU-based variants.
Finally, we show in Figure 5 an example of a diverging FP8DPA run, comparing it with the successful FOG-max training. This emphasises the importance of tracking tensor-level metrics such as kurtosis to potentially predict later divergences, before common global metrics like the loss and gradient norms show any symptoms of divergence. In this example, while loss irrefutably diverged around the 15B token mark and the gradient norm consistently spiked no earlier than 12B tokens, the QKV kurtosis was already diverging from the expected sub-linear growth consistently seen across different architectures as early as the 3B mark, giving a potential early divergence sign.
Figure 5: Training dynamics of a failed and a successful FP8DPA run. Kurtosis exhibits atypical behaviour much earlier than when the loss diverged.
@@ 5 Experimental Results
We perform extensive experiments to verify our architecture across several scales. We use the FineWeb-Edu [21] text corpus, filtering out any web opt-out domains with robots.txt, resulting in a rigorous data-compliant corpus [5]. The data is tokenised using a 131K vocabulary BPE tokenizer. We keep a consistent context length of 4096 during all experiments. In terms of the optimisation algorithm, we use AdamW [17] with default hyperparameters. Our learning rate schedule is comprised of three phases: Warm-up, Steady, and Decay phases (WSD), as it has been shown to provide equivalent performance to the cosine schedule [8], while allowing to train beyond fixed training durations. For the models, we train 390M, 1.5B and 8B parameter models for different token counts, specified at each experiment. Our baseline architecture follows the Llama3 8B model design [7], with the 390M and 1.5B being adapted to their respective sizes. Since Llama3 uses a gated linear unit, unlike the OP and FOG variants, we increase the FFN sizes of OP and FOG to maintain an equal parameter count. Further details regarding architectures and hyperparameters are available in the Appendix.
Our hardware infrastructure consists of nodes with 4 Nvidia Grace Hopper GPUs each. Our distributed training framework is adapted from Megatron-LM [27], which uses Transformer Engine [2] FP8 recipes. With 390M parameters, our experiments reach 50B tokens. We scaled 1.5B experiments to 125B tokens to obtain more meaningful evaluations. In addition to the absence of late-in-training outlier amplification from FOG’s non-gated activation functions and our kurtosis progression guarantees, we further validate our method’s stability on long data regimes by continuing pretraining FOG-max up to 420B tokens. Finally, we scale the model size to 8B and train for 20B tokens. We show the divergence of other architectures with FP8DPA while FOG variants converge and match the baseline Llama3 BF16 loss, while being 35-40% faster. During all experiments, we use the delayed scaling strategy, with a margin of zero and a history length of 1024 steps.
We make our implementation, along with reproducibility steps for our experiments, public under the anonymized repository https://github.com/anonymous4375934/FOG.
@@ 5.1 FP8 stability
Figure 6: Cross-entropy loss plots across different architectures with FP8DPA training. No other tested architecture was able to surpass the 20B token mark without diverging.
Another interesting observation from these experiments is the tendency of larger models to diverge in later stages of training compared with similar but smaller models. We validate its consistency across architectures, as presented in Table 2. This observation has not been raised before, possibly due to the longer time needed for FP8-only models to diverge, in contrast to FP8DPA training. While this trend could have many practical implications, exploring it fully falls outside the scope of this work, and we encourage future research in this direction.

### v2

@@ 1 Introduction
Thus, the development of efficient training techniques has become increasingly essential. One of the main research avenues for efficiency is the use of lower-precision number formats to accelerate training on appropriate hardware accelerators. Recently, the use of 8-bit floating-point (FP8) formats has shown promising results [3, 6, 24]. However, the widespread adoption of current approaches is still limited due to suboptimal throughput benefits. One cause of slowdowns is the use of higher precision in those General Matrix Multiplications (GEMMs) which are most sensitive, such as attention score computation, while another issue is the overhead caused by more granular FP8 scaling strategies. One of the key challenges in 8-bit LLM training originates from the relatively narrow dynamic range offered by FP8 formats and thus higher risk of underflows and overflows, especially with the prevalence of large outlier features observed in the LLM’s neural activations during training [4, 30, 9]. We formalize the effect of outliers on quantization later in Appendix C. To mitigate this issue, modern FP8 training recipes utilise various scaling techniques before casting from higher-precision formats—typically BF16 [14] for activations—to FP8 formats used in matrix multiplications. These scaling approaches help maximize the effective use of FP8’s limited dynamic range, reducing the risk of underflows and overflows.
@@
Recent work has introduced promising FP8 training recipes by employing multiple scaling factors per single tensor [3], allowing for a finer and more precise casting to lower precision. Yet, this comes with an efficiency overhead, diminishing the large gains initially expected from using FP8. Another strategy involves adjusting the standard SwiGLU-based transformer architecture [28] to prevent emergent outliers from occurring [6]. This area of optimization remains underexplored, as most works focus on FP8 GEMMs within the linear projections within the transformer, while maintaining higher precision for other GEMMs, namely those involved in the dot product attention mechanism. We refer to such training strategies simply as FP8 training. We label the approach of also including FP8 attention computation as FP8DPA.
@@
In this paper, we introduce FOG: the Fast and Outlier-Guarded set of LLM architectures specifically designed to mitigate large activation outliers and enable efficient large-scale FP8 training with low-overhead scaling strategies. For the first time to our knowledge, this approach enables FP8 GEMMs not only in the linear projection, but also within the attention mechanism of each transformer block, achieving unprecedented throughput improvements of up to 43% in the 8B parameter model scale, while maintaining equivalent downstream performance compared to higher precision baselines. In addition, we present a comprehensive recipe for monitoring, explaining, and predicting training instabilities that might not surface in the early stages of training. This approach provides researchers with greater confidence in the long-term stability of FP8 training recipes, reducing the need for costly, full-scale experiments when using new architectures. Furthermore, we provide an interestingly useful observation about larger models’ tendency to diverge later in training with FP8.
@@
  • We introduce the FOG set of architectures, designed to minimise outlier features during training. Our recipe allows stable training with FP8 computation of all GEMMs inside the transformer blocks, surpassing the throughput of the standard BF16 approach by up to 43%.
  • Our design achieves equivalent quality results to BF16 baselines, while providing a significant speed-up. We empirically attest both performance and stability on various model sizes (0.4B, 1.5B, 8B) and data regimes up to 15x the Chinchilla optimal data budget [12].
  • We show the flexibility of FOG design as it can be adapted to several architectures including various families of activation functions and even Mixture-of-Experts (MoE) settings.
@@ 2 Background
Further examination has shown that not only is SwiGLU an outlier amplifier, but Gated Linear Units (GLUs) in general, as well as pre-normalization layers, suggesting that improper signal propagation is the root cause of outliers [9]. Removing these components and equipping transformers with QK entropy regularization mechanisms such as QK RMS Normalization [10], producing the Outlier Protected (OP) architectures [9], has been shown to diminish late-stage outliers observed by orders of magnitude, while providing equivalent prediction quality. While OP architectures were shown to be beneficial for post-training quantization, its use for FP8 pre-training remains unexplored. Finally, an alternative to pre-normalization layers are post-normalization layers [18]. Long data regime trainings have confirmed their superiority in terms of training stability with the standard BF16 mixed precision training [21].
@@ 3 FOG: Fast and Outlier-Guarded FP8-suited architectures
Further, the input of the first transformer block is scaled by $\sigma^{-1}$ to maintain unit variance activations at initialization, where $\sigma$ is the chosen standard deviation of the network’s random initialization. Finally, to enhance performance, a learnable normalization block is applied before the residual connections. This takes the form of a LayerScale [31] block, $\mathrm{LayerScale}_{\boldsymbol{\gamma}}(\mathbf{x}):=\boldsymbol{\gamma}\odot\mathbf{x}$, where $\boldsymbol{\gamma}\in\mathbb{R}^{D}$ is a learnable gain vector, or an RMSNorm block, resulting in a post-normalized architecture [18]. In both cases, the learnable gains vector is initialized to $\mathbf{1}/\sqrt{\mathrm{num\_layers}}$ and keeps the residual branch unnormalized, allowing proper signal propagation [9]. Our architecture suite is specified in Table 1, and further details are available in Appendix B.
@@
While the OP architecture already offers several guards to prevent large outliers from occurring, we observed that it remains an impractical choice for FP8DPA training. In Section 5.1 we show that, like all other architectures tested, it suffers a fatal loss divergence early during training. We isolate the two components responsible for OP’s incompatibility with FP8DPA training: the trainable QK RMSNorm gains vector $\boldsymbol{\gamma}$, and the lack of any normalization. We identify the use of post-normalization as not prone to the outlier tendency pre-normalization networks have.
Figure 2: From OP to FOG-opt step by step. Comparison of 390M models under FP8DPA training. The first architecture to diverge is OP, while OP with frozen QK RMSNorm gains survives the stable phase of training. It still, however, experiences a significant divergence during the learning rate cooldown, which starts around 42B tokens in. The architecture that converges, FOG-opt, is the result of adding post-normalization to the previous recipe. Gradient norm reported is the 200-rolling-window mean and 5%-95% quantile bands.
Figure 2 ablates the components transitioning from OP to FOG-opt. We can see that freezing the trainable QK RMSNorm gains results in a significantly more stable training. We attribute the early divergence of OP to the fact that uncontrolled QK normalization leads to an explosion of its gains when training in low precision. Note that these gains are generally not weight-decayed. We experimentally observe this explosion, confirm that using $L_{2}$ regularization helps delay the divergence. We finally opted for freezing the gains to a constant value as it is simpler and sufficient, doesn’t compromise performance, and offers a small speedup. Our ablations highlight that a constant value for the gains slightly greater than 1 improves loss. Therefore, to retain its benefit after removing the $\boldsymbol{\gamma}$ gains vector, we increase the standard $s=1/\sqrt{D_{qk}}$ attention softmax scale–a tiny optimization trick offering equivalent attention score matrix $\mathbf{S}$:
@@
Prior works also favored post-normalization over pre-normalization [21], providing evidence of their better stability in BF16 training. We extend this observation to our FP8 setting and we confirm that learnable LayerScale blocks alone, even with controlled QK regularization, cannot ensure convergence during this last phase. We attribute this late divergence of OP to the fact that LayerScale blocks without normalizations are not enough to handle FP8 outliers, potentially due to the considerable changes in model statistics following the learning rate decay, that are summed up in the residual connections resulting in huge activation outliers for last layers, as highlighted by the increasing pattern of outliers on each transformer block’s output in Figure 3. We note that we initially tested the idea of cooling down the previously constant weight decay during the learning cooldown phase, aiming to conserve model weights’ norm [16]. The Appendix G shows that such intervention has no noticeable effect on stability neither performance. We chose to conserve this decision for all FOG runs for consistency and fair comparison across ablations.
@@ 4 Long-term outlier dynamics
Figure 4: Loss and kurtosis training dynamics of 1.5B FOG-max and Llama3 models trained for over 100B tokens with BF16 precision. Loss reported is the 200-rolling-window mean and 5%-95% quantile bands.
@@
Activation Functions
Baseline Llama exhibits late divergence during FP8 training (with attention in BF16), which has been attributed solely to the quadratic behavior of its gated activation function—emerging when weights become sufficiently aligned late in training [6]. In our extended 450B token run using the FOG-max architecture, we employ the inherently quadratic xIELU activation function, see Equation (2), and observe stable training with kurtosis levels orders of magnitude lower than those of baseline Llama. In fact, modifying the FOG-max architecture to use the SwiGLU activation function resulted in stable FP8DPA training behaviour, as disscussed in Appendix D. These results strongly suggests that architectures biased towards low kurtosis activations during training enable the stable use of quadratic activations, and challenges the completeness of prior explanations. This is particularly interesting given that such activations are known to produce linear gradients, which benefit the backward pass—likely contributing to FOG-max’s superior performance over GeLU-based variants as seen in Section 5.3.
Long-term outlier growth
These architectures exhibit a sub-linear to logarithmic trend in the long-term growth of QKV outliers, as consistently shown by kurtosis in Figure 4. This behavior supports their robustness to FP8DPA, as it suggests that prohibitively longer training would be required to see a substantial increase in kurtosis. Our extended run is consistent with the hypothesis as it does not exhibit any sign of divergence.
Kurtosis early signal
Figure 5 shows an example of a diverging FP8DPA run, comparing it with the successful FOG-max training. This emphasises the importance of tracking tensor-level metrics such as kurtosis to potentially predict later divergences, before common global metrics like the loss and gradient norms show any symptoms of divergence. In this example, while loss irrefutably diverged around the 15B token mark and the gradient norm consistently spiked no earlier than 12B tokens, the QKV kurtosis was already diverging from the expected sub-linear growth consistently seen across different architectures as early as the 3B mark, giving a potential early divergence sign.
Figure 5: Training dynamics of a failed and a successful FP8DPA run. Kurtosis exhibits atypical behaviour much earlier than when the loss diverged. Gradient norm reported is the 200-rolling-window mean and 5%-95% quantile bands.
@@ 5 Experimental Results
We perform extensive experiments to verify our architecture across several scales. We use the FineWeb-Edu [23] text corpus, filtering out any web opt-out domains with robots.txt, resulting in a rigorous data-compliant corpus [5]. The data is tokenised using a 131K vocabulary BPE tokenizer. We keep a consistent context length of 4096 during all main experiments. In terms of the optimisation algorithm, we use AdamW [19] with default hyperparameters. Our learning rate schedule is comprised of three phases: Warm-up, Steady, and Decay phases (WSD), as it has been shown to provide equivalent performance to the cosine schedule [8], while allowing to train beyond fixed training durations. For the models, we train 390M, 1.5B and 8B parameter models for different token counts, specified at each experiment. Our baseline architecture follows the Llama3 8B model design [7], with the 390M and 1.5B being adapted to their respective sizes. Since Llama3 uses a gated linear unit, unlike the OP and FOG variants, we increase the FFN sizes of OP and FOG to maintain an equal parameter count. Further details regarding architectures and hyperparameters are available in Appendices A and B.
Our hardware infrastructure consists of nodes with 4 Nvidia Grace Hopper GPUs each. Our distributed training framework is adapted from Megatron-LM [29], which uses Transformer Engine [2] FP8 recipes. With 390M parameters, our experiments reach 50B tokens. We scaled 1.5B experiments to 125B tokens to obtain more meaningful evaluations. In addition to the absence of late-in-training outlier amplification from FOG’s non-gated activation functions and our kurtosis progression guarantees, we further validate our method’s stability on long data regimes by continuing pretraining FOG-max up to 450B tokens. Finally, we scale the model size to 8B and train for 20B tokens. We show the divergence of other architectures with FP8DPA while FOG variants converge and match the baseline Llama3 BF16 loss, while being 35-43% faster. During all experiments, we use the FP8 delayed scaling strategy, with a margin of zero and a history length of 1024 steps.
We make our implementation, along with detailed steps for our experiments, public under the repository https://github.com/anonymous4375934/FOG.
@@ 5.1 FP8 stability
Figure 6: Cross-entropy loss plots of different architectures with FP8DPA training. No other tested architecture was able to surpass the 20B token mark without diverging at any scale.
Another interesting observation from these experiments is the tendency of larger models to diverge in later stages of training compared with similar but smaller models. We validate its consistency across architectures, as presented in Table 2. This observation has not been raised before, possibly due to the longer time needed for FP8 settings (with BF16 attention) to diverge, in contrast to FP8DPA training. While this trend could have many practical implications, exploring it fully falls outside the scope of this work, and we encourage future research in this direction.
@@
5.2 Efficiency
Standard context length
Table 3 explores the efficiency of FOG at 1.5B and 8B model scales under FP8DPA training, using a standard context length of $4096$. We compare our set-up with the BF16 baseline and the stable Llama FP8 training with SmoothSwiGLU, which is, to the best of our knowledge, the only dense architecture proposal demonstrated to work at scale with FP8. Note that the SmoothSwiGLU cannot benefit from enabling FP8 GEMMs in the attention mechanism, as it was shown to suffer a big loss divergence in Figure 6. Note the increase in throughput gains as the model size increases. The GEMM input tensors increase in size and consume significantly more time during the overall forward-backward pass, compared with other operators.
Size Model Precision
Throughput
(tokens/second/GPU)
8B Llama BF16 9105
Llama+SmoothSwiGLU FP8 12228 (+34.3%)
FOG-max FP8DPA 12344 (+35.5%)
FOG-opt FP8DPA 12414 (+36.3%)
FOG-flash FP8DPA 12764 (+40.2%)
1.5B Llama BF16 46470
FOG-max FP8DPA 53551 (+15.2%)
FOG-opt FP8DPA 53877 (+15.9%)
FOG-flash FP8DPA 54848 (+18.0%)
Llama+SmoothSwiGLU FP8 54903 (+18.1%)
Table 3: Training throughput measures with FOG versus other baselines. Using eight GH200 nodes with Zero-1 sharding [26] for 8B models and a single GH200 node for 1.5B models. Notably, in the 8B scale, all FOG variants outperform other architectures.
Long-context scenario
Enabling stable training under FP8DPA regime leads to great benefits in long context scenarios, because the throughput becomes bottlenecked by the dot product attention computation with quadratic complexity. As a result, we observe larger speed-up gaps between FOG and prior FP8 approaches as sequence length increases, as demonstrated in Table 4.
Context FOG-flash Llama+SmoothSwiGLU     Speed-up Gap
4096 (TP=1) +42.6% +38.5%     +4.1%
8192 (TP=1) +43.5% Out Of Memory
8192 (TP=2) +39.1% +34.2%     +4.9%
16384 (TP=2) +38.8% +31.1%     +7.7%
Table 4: Training throughput gains under varying sequence lengths (relative to Llama BF16), performed at 8B scale using 8 GH200 nodes with a global batch size of 1024. Increasing to longer contexts required enabling Tensor Parallelism (TP). Raw throughput values reported in Table 11.